# 0Year 0Year 0Year Optimal Inference of Asynchronous Boolean Network Models

### v1

@@ Abstract
Associations between phenotype and genomic and epigenomic markers are often derived by correlation. Systems Biology aims to make more robust connections and uncover broader insights by modeling the cellular mechanisms that produce a phenotype. The question of choosing the modeling methodology is of central importance. A model that does not capture biological reality closely enough will not explain the system’s behavior. At the same time, highly detailed models suffer from computational limitations and are likely to overfit the data. Boolean networks strike a balance between complexity and descriptiveness and thus have received increasing interest. We previously described an algorithm for fitting Boolean networks to high-throughout experimental data that finds the optimal network with respect to the information in a given dataset. In this work, we describe a simple extension that enables the modeling of asynchronous dynamics, i.e. different reaction times for different network nodes. Our approach greatly simplifies the construction of Boolean network models for time-series datasets, where asynchronicty often occurs. We demonstrate our methodology by integrating real data from three datasets, and provide an implementation that can be used by the community for network reconstruction using any high-throughout dataset. Our approach significantly expands the applicability of the Boolean network model to experimental data.
@@
Cellular processes that alter the functional profile of the cell are subject to complex, adaptable regulation. This regulation is achieved by networks of interactions between molecules that integrate environmental signals and the current state of the cell (Arenas et al., 2015; Hackett et al., 2016; Eichenberger et al., 2004). In order to understand these networks, various modeling methodologies have been proposed, varying in their degree of complexity and expressiveness(Karlebach and Shamir, 2008). Due to limited knowledge of the reactions involved in carrying out each regulatory interaction, Boolean models, first introduced by Kauffman (Kauffman, 1969), are frequently chosen as a general-purpose modeling methodology (Fox et al., 2024; Samaga et al., 2009; Schwab et al., 2017). Several questions arise when reconstructing a Boolean network from data. First, as with any computational task, one must ask whether the process can be accomplished efficientlyKarlebach2012. The second question is how to find the correct trade-off between model size and the fit to the data(Rissanen, 1983; Akaike, 1998; Schwarz, 1978). Finally, a method to map continuous or discrete measurements into binary values is needed (Hopfensitz et al., 2012; Glaz et al., 2001; Shmulevich and Zhang, 2002; Berestovsky and Nakhleh, 2013). In addressing the first question, Karlebach and Shamir showed that the problem of optimally fitting a Boolean model to binary data is NP-Complete (Karlebach and Shamir, 2008), making a solution impractical in the worst case and likely challenging for many other instances. The second question is usually addressed by independently minimizing the fit of individual regulatory logic tables to the binarized values of their targets (Margolin et al., 2006; Lahdesmaki, 2003; Faith et al., 2007). Bayesian approaches have also been proposed, but these arguably deviate from the simplicity of the Boolean model by introducing a probabilistic parameters (Gat-Viks et al., 2004). Karlebach and Robinson proposed a non-probabilistic criterion that is based on Kolmogorov Complexity (Kolmogorov, 1968b) ,and takes into account both the fit to the data and the total size of the model (Karlebach and Robinson, 2023a), and proposed an algorithm for optimizing according to it. As for the third question, while a gold-standard has not been agreed upon, various methodologies have been developed that generate useful mappings in practice (Berestovsky and Nakhleh, 2013). An often overlooked challenge in modeling is accounting for variation in reaction times. A gene may be activated at a different rate than another gene that has a regulatory association to it, and this can result in a time-series dataset where not all regulations take effect at every sampling point. A related issue occurs when consecutive sampling times capture the same network state repeatedly. Since the data is usually affected by noise, it may be hard to know when the exact same state has been sampled twice. To address these issues, our goal in this paper is to extend the methodology described in (Karlebach and Robinson, 2023a) for Boolean networks with asynchronous dynamics, thereby obtaining a rigorous reconstruction methodology that can handle time-series data where the sampling rate has not been optimized, or where the rates of different regulatory interactions differ significantly. In the following section, we specify and extend our methodology and explain its application to asynchronous network modeling. We also show how heuristics that address the computational complexity of the problem can be used in the case of asynchronous dynamics. In the Results section we describe modeling of real data with asynchronous Boolean networks. In the Conclusion section we summarize our findings and assess the impact of the method on future modeling efforts.
@@ 3 Results
To test our method on real experimental data, we obtained the microarray dataset GSE49650 of synchronized yeast cells from the Gene Expression Omnibus. Preprocessing of the .CEL files was done using the rma function in the Bioconductor package affy, using default argument values. Additionally, every trajectory (time-series) of every gene was smoothed using the R functions smooth.spline with smooth parameter 0.5 and then approxfun (in the stats package) with default argument values(Becker et al., 1988). The x-coordinate values for the smoothing were the times at which the measurements were taken in minutes, and the y-coordinate values were the array intensities. We used the BASCA method as implemented in the R package Binarize (Hopfensitz et al., 2012), with default arguments, for mapping from continuous to Boolean values. Every trajectory was binarized separately. We used the yeast cell cycle model from Cho et al. (Kolmogorov, 1968a), with the edges as the candidate regulatory connections. Complexes were modeled using the expression of one of their genes. After fitting the model, the percentage of mismatches was 16.64 %.

### v2

@@ Abstract
Associations between phenotype and genomic and epigenomic markers are often derived by correlation. Systems Biology aims to make more robust connections and uncover broader insights by modeling the cellular mechanisms that produce a phenotype. The question of choosing the modeling methodology is of central importance. A model that does not capture biological reality closely enough will not explain the system’s behavior. At the same time, highly detailed models suffer from computational limitations and are likely to overfit the data. Boolean networks strike a balance between complexity and descriptiveness and thus have received increasing interest. We previously described an algorithm for fitting Boolean networks to high-throughout experimental data that finds the optimal network with respect to the information in a given dataset. In this work, we describe a simple extension that enables the modeling of asynchronous dynamics, i.e. different reaction times for different network nodes. Our approach greatly simplifies the construction of Boolean network models for time-series datasets, where asynchronicty often occurs. We demonstrate our methodology by integrating real data from transcriptomics experiments, and provide an implementation that can be used by the community for network reconstruction using any high-throughout dataset. Our approach significantly expands the applicability of the Boolean network model to experimental data.
@@
Cellular processes that alter the functional profile of the cell are subject to complex, adaptable regulation. This regulation is achieved by networks of interactions between molecules that integrate environmental signals and the current state of the cell (Arenas et al., 2015; Hackett et al., 2016; Eichenberger et al., 2004). In order to understand these networks, various modeling methodologies have been proposed, varying in their degree of complexity and expressiveness(Karlebach and Shamir, 2008). Due to limited knowledge of the reactions involved in carrying out each regulatory interaction, Boolean models, first introduced by Kauffman (Kauffman, 1969), are frequently chosen as a general-purpose modeling methodology (Fox et al., 2024; Samaga et al., 2009; Schwab et al., 2017). Several questions arise when reconstructing a Boolean network from data. First, as with any computational task, one must ask whether the process can be accomplished efficientlyKarlebach2012. The second question is how to find the correct trade-off between model size and the fit to the data(Rissanen, 1983; Akaike, 1998; Schwarz, 1978). Finally, a method to map continuous or discrete measurements into binary values is needed (Hopfensitz et al., 2012; Glaz et al., 2001; Shmulevich and Zhang, 2002; Berestovsky and Nakhleh, 2013). In addressing the first question, Karlebach and Shamir showed that the problem of optimally fitting a Boolean model to binary data is NP-Complete (Karlebach and Shamir, 2008), making a solution impractical in the worst case and likely challenging for many other instances. The second question is usually addressed by independently minimizing the fit of individual regulatory logic tables to the binarized values of their targets (Margolin et al., 2006; Lahdesmaki, 2003; Faith et al., 2007). Bayesian approaches have also been proposed, but these arguably deviate from the simplicity of the Boolean model by introducing a probabilistic parameters (Gat-Viks et al., 2004). Karlebach and Robinson proposed a non-probabilistic criterion that is based on Kolmogorov Complexity (Kolmogorov, 1968) ,and takes into account both the fit to the data and the total size of the model (Karlebach and Robinson, 2023a), and proposed an algorithm for optimizing according to it. As for the third question, while a gold-standard has not been agreed upon, various methodologies have been developed that generate useful mappings in practice (Berestovsky and Nakhleh, 2013). An often overlooked challenge in modeling is accounting for variation in reaction times. A gene may be activated at a different rate than another gene that has a regulatory association to it, and this can result in a time-series dataset where not all regulations take effect at every sampling point. A related issue occurs when consecutive sampling times capture the same network state repeatedly. Since the data is usually affected by noise, it may be hard to know when the exact same state has been sampled twice. To address these issues, our goal in this paper is to extend the methodology described in (Karlebach and Robinson, 2023a) for Boolean networks with asynchronous dynamics, thereby obtaining a rigorous reconstruction methodology that can handle time-series data where the sampling rate has not been optimized, or where the rates of different regulatory interactions differ significantly. In the following section, we specify and extend our methodology and explain its application to asynchronous network modeling. We also show how heuristics that address the computational complexity of the problem can be used in the case of asynchronous dynamics. In the Results section we describe modeling of real data with asynchronous Boolean networks. In the Conclusion section we summarize our findings and assess the impact of the method on future modeling efforts.
@@ 3 Results
To test our method on real experimental data, we obtained the microarray dataset GSE49650 of synchronized yeast cells from the Gene Expression Omnibus. Preprocessing of the .CEL files was done using the rma function in the Bioconductor package affy, using default argument values. Additionally, every trajectory (time-series) of every gene was smoothed using the R functions smooth.spline with smooth parameter 0.5 and then approxfun (in the stats package) with default argument values(Becker et al., 1988). The x-coordinate values for the smoothing were the times at which the measurements were taken in minutes, and the y-coordinate values were the array intensities. We used the BASCA method as implemented in the R package Binarize (Hopfensitz et al., 2012), with default arguments, for mapping from continuous to Boolean values. Every trajectory was binarized separately. We used the yeast cell cycle model from Cho et al. (Cho et al., 2019), with the edges as the candidate regulatory connections. Complexes were modeled using the expression of one of their genes. After fitting the model, the percentage of mismatches was 16.64 %.